OpenStudy (anonymous):

through (1, -5) and (a) parallel, (b) perpendicular to the line through (3,5),(3, -7).

OpenStudy (anonymous):

Please show some of your own work. You MUST have something to show us so we can help you more specifically.

OpenStudy (anonymous):

w8.. ill try to show you

OpenStudy (anonymous):

// to L1 Eq'n of L2 mL1 = y2-y1/x2-x1 mL1 = -7-5 / 3-3 = - infinity <---- this part i got stock

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@OpiGeode